Can someone help me with these problems, thank you!
3. ( points). A circular coil of 4 loops with radius R 50 mm moves perpendicularly into a magnetic field B which increases from 0.0?to 18.0 T in 3.0 s If the resistance of the coil is R 2.0 ?, Bttal 10.01 RMN (a) what is the magnitude of the induced current, Din the cail? I- (Ab)k 50 10 m -|174,10? 1-134 10 A (b) what is the direction of the current I in the coil seen face on (counter clockwise CCW or clockwise CW) if the direction of the magnetic field Bi created by the current is as shown? Direction: 4. (25 points). Calculate the emre induced if a right triangle (of height h-0.30 m and base b = 0.15 m) is moved at an angle ?-45° into to a uniform magnetic field initially of strength B-2.5 T which then decreases to B = 1.5 T in 2.0 s. no 100 ps 450 h = 0.30 m b=0.15m atv 203 2.0
Answer #1

solving first question

Q3

induced emf =change in flux/time

change in flux=0.1413

induced emf=0.1413/3=0.047 V

induced current=emf/R=0.0235A=23.6 mA

(b) Induced current will be such that it will oppose the increase of magnetic field (By Lenz's law)

so direction is clockwise if seen from right side of coil